# A Survey on Bundle Recommendation: Methods, Applications, and Challenges

### v1

@@ 3.4. Summary of Existing Discriminative BR Research
(Deng et al., 2020; Chang et al., 2020; Zhao et al., 2022; Wei et al., 2023; Wang et al., 2023c; Ma et al., 2022)
(Liu et al., 2023b; Zhu et al., 2023c; Ke et al., 2023b, a; Ren et al., 2023; Zhang et al., 2022)
@@
(Li et al., 2023c; Lu et al., 2022; Vijaikumar et al., 2021; Zhang et al., 2021b; Liu et al., 2023a; Yu et al., 2022)
- (Chen et al., 2019b; Sun et al., 2022a; Wang et al., 2021b)
Category, Attributes, Genre - (Xu et al., 2023; Yang et al., 2023; Li et al., 2023b, a)
@@ 4.1. Generative BR Task
(24) $\phi^{*}=\arg\max_{\phi}\sum_{u\in\mathcal{U}_{\text{train}}}\log p(g_{\text{gen}}(u,b,\mathbf{X}_{M\times N},\mathbf{Y}_{M\times O},\mathbf{Z}_{N\times O},c;\phi)\mid u),$
@@ 4.2. Representation Learning from Item Level
Category Summarization References
Embedding Models MF-based (Yang et al., 2024b; Zhang et al., 2024)
DNN-based(RNN, LSTM et al.)
(Bai et al., 2019; Wei et al., 2022; Han et al., 2017; Chen et al., 2019a; Li et al., 2017; Deng et al., 2023b; Wang et al., 2023a)
(Zheng et al., 2023; Gu et al., 2021; Liu et al., 2023c; Liu and Cao, 2020; Deng et al., 2021, 2023a; Zhou et al., 2019)
(Ma et al., 2024c, b; Xu et al., 2024; Zhang et al., 2024; Liu et al., 2024; Gao et al., 2023b)
GNN-based(GCN, GAT et al.)
@@
(Zheng et al., 2023; Qi et al., 2021; Deng et al., 2021; Li et al., 2023d)
(Ma et al., 2024c; Ding et al., 2023; Yu et al., 2023; Zhu et al., 2023b)
Feature Engineering Feature Extraction(cost, time, price et al.)
(Yu et al., 2015; Fang et al., 2014; Chen et al., 2015; Xie et al., 2010)
(Benouaret and Lenne, 2017, 2016; Banerjee et al., 2020; Chen et al., 2016)
CF-based(UserCF, ItemCF)
@@ 4.2.1. Embedding Models
Similar with representation learning of Discriminative BR, both DNN-based and GNN-based models are frequently utilized in Generative BR to learn representations in many studies (Chang et al., 2021; Vijaikumar et al., 2021; Zheng et al., 2023; Yang et al., 2024b; Bai et al., 2019; Chen et al., 2019a), etc.
@@
MF-based. BundleNAT (Yang et al., 2024b) introduces two kinds of signal to learn the needed dependency information, i.e., the preference signal and the item compatibility signal. It employs MF to extract the preference signal from user-item interactions and uses GNNs to capture item compatibility signal from the co-occurrence graph, which can strengthen the intra-relatedness within the bundle. MatTrip (Zhang et al., 2024) utilizes Funk-SVD for feature completion to extract and complete the user’s interest.
@@
DNN-based. BGN (Bai et al., 2019) uses CNN as an encoder to extract user context information, embedding user based on user purchase history. Conna (Wei et al., 2022) adopts self-attention to learn the representations for mixture types of candidate objects. Users’ interests are usually distributed across multiple aspects and a fine-grained representation learning of user interests from multiple aspects is conducted by (Deng et al., 2023a). However, this approach does not fully consider the comprehensiveness and dynamism of user interests. To address this limitation, a multi-view multi-aspect neural network representation learning method is introduced by (Deng et al., 2023b), which can more accurately model the multi-aspect interests of users and enhance the accuracy of next-basket recommendations. In works such as (Chen et al., 2019a; Han et al., 2017; Li et al., 2017; Ma et al., 2024c, b), items are represented using a multi-modal embedding model based on multimodal inputs, including image, text, category of the items. Recent advancements in large-scale multimodal foundation models, such as BLIP (Li et al., 2022b) and CLAP (Wu et al., 2023a), have paved the way for studies like CLHE (Ma et al., 2024c) and Bundle-LLM (Liu et al., 2024) to harness these models for extracting the multimodal features of items. These studies have introduced self-attention mechanism to fuse multimodal information and multi-item representation. Furthermore, Bundle-LLM (Liu et al., 2024) employs CF method to obtain relational features from user-item interactions and bundle-item affiliations, which can enrich the representation of items. CIRP (Ma et al., 2024b) learns high-quality representations of item with multimodal encoders and pre-trained targets including image-text contrastive loss and cross-item contrastive loss. In order to capture the compatibility of items, a masked item prediction task based on a bidirectional Transformer encoder architecture is devised by (Chen et al., 2019a) to capture the item interactions in a bundle. Bi-LSTM is utilized in (Han et al., 2017) to learn item compatibility within a bundle. Similarly, MatTrip (Zhang et al., 2024) processes user interest and geographic information separately by two independent Bi-LSTM encoders. In the domain of service package generation, the process begins with parsing textual descriptions of services and mashups to distill semantic features. This is achieved through techniques such as Latent Dirichlet Allocation (LDA) and Natural Language Processing (NLP) methods, which convert text into vector representations that capture essential topics and semantics, as discussed in the works (Gu et al., 2021) and (Liu et al., 2023c). Services and mashups are initially represented as static embeddings, using pre-trained word embeddings like Word2Vec or GloVe to encapsulate their inherent attributes and user requirements (Liu and Cao, 2020; Qi et al., 2021).
@@
GNN-based. To address the dynamic nature of services, a dynamic graph neural network is constructed to model the evolving relationships between services. The graph typically includes nodes for services, mashups, categories, and sometime words, with edges defined by various relationships such as co-occurrence, usage frequency, and semantic associations (Liu et al., 2023c). Graph convolution operations, potentially involving GraphTransformer or Graph Convolutional Networks (GCNs), are applied to update the static embeddings by aggregating information from neighboring nodes in the graph. This allows for the dynamic adjustment of service representations in accordance with the network’s topology (Liu and Cao, 2020; Qi et al., 2021). The graph attention mechanism is leveraged in (Vijaikumar et al., 2021) to capture the high-order relationship between users, items and bundles, which can understand the constitution of a bundle better. BGGN (Chang et al., 2021) learns user and item embeddings through GNNs, and uses GCN to learn the complex interactions upon re-constructed bundle graph. Hypergraph convolution operations are performed on hypergraph to learn better and intent-aware representations of hyperedges (Yu et al., 2023; Li et al., 2023d). TOG (Ding et al., 2023) first defines category combinations in outfit as templates, building two bipartite graphs (user-template and user-item), and then employ GNNs to obtain the enhanced representations of users, templates and items. In drug package scenario, DPG (Zheng et al., 2023) models item relationships by constructing item interaction graph based on medical records and domain knowledge. It devises a message-passing neural network to capture drug interaction (synergism, antagonism, etc.) and uses DNNs to obtain the patient embeddings.
@@ 4.2.2. Feature Engineering
In the context of personalized travel package recommendations, Yu et al. leverage CF to dynamically extract user preferences from location-based social network data, constructing user profiles that reflect individual preferences for various points of interest (POIs) (Yu et al., 2015). Similarly, Fang et al. propose a framework that integrates both POIs and travel packages, employing a score inference model to personalize trip recommendations, which takes into account user-based preferences and temporal properties, inferred through CF techniques (Fang et al., 2014). Benouaret et al. further extend the application of CF by designing a package recommendation framework for trip planning. Their system emphasizes the importance of diversity and relevance in package recommendations, using a composite recommendation system inspired by CF to group diverse POIs into packages that align with user preferences and budget constraints (Benouaret and Lenne, 2016). Building on this, Benouaret et al. aim to recommend top-k packages to users, with each package constituted by a set of POIs under specified constraints. They introduce a scoring function that considers user preferences, diversity, and popularity, all of which are integral components of CF-based representation learning (Benouaret and Lenne, 2017).
@@ 4.3. Bundle Generation
Category Techniques References
@@
Beam Search (Bai et al., 2019; Chang et al., 2021; Yu et al., 2015; Zhang et al., 2024; Sun et al., 2023b)
Breadth-First Search (Liu and Cao, 2020)
@@
(Zheng et al., 2023; Bai et al., 2019; Han et al., 2017; Zhou et al., 2019; Gao et al., 2023b)
(Li et al., 2017; Yu et al., 2023; Zhu et al., 2023b; Zhou et al., 2022)
@@
(Wei et al., 2022; Yang et al., 2024b; Chen et al., 2019a)
(Sun et al., 2023b; Wang et al., 2023a)
@@ 4.3.1. Search-based
Greedy search and beam search have also been employed to construct personalized travel routes (Gu et al., 2020; Fang et al., 2014; Yu et al., 2015; Zhang et al., 2024). TRAR (Gu et al., 2020) uses greedy search to generate candidate set by gradually adding POIs to build the optimal travel package that satisfy multiple user-specific constraints. Yu et al. (Yu et al., 2015) utilize a path-tree based beam search method, focusing on iteratively selecting POIs based on the highest transition probabilities to form travel routes. In contrast, the work by Zhang et al. (Zhang et al., 2024) integrates user preferences, geographic information, and a grid beam search approach within an iterative process to select candidate POIs, generating travel routes that meet individualized needs and constraints. Furthermore, dynamic programming and integer programming have been used to solve search problem of forming optimal travel packages (Chen et al., 2016; Lim et al., 2015).
@@ 4.3.2. Model-based
Search-based methods generate bundles from the candidate set by searching for the optimal combinations. These methods can explore different combinations more intuitively, but face two problems: first, the search process often relies on predefined heuristics, so the generation effect may not be flexible enough in complex scenarios. Second, the search space expands drastically with the increase of the candidate set, leading to a significant increase in computational overhead. At this point, model-based can cope with these problems and better capture complex relationships, generating bundles with higher quality. As introduced in Section 4.2, embedding models and feature extraction techniques are used for distilling features, representing data, and learning compatibility between items. The subsequent phase involves the utilization of these encoded representations by generative models, including but not limited to Recurrent Neural Networks (RNNs) and Generative Adversarial Networks (GANs). These models are tasked with the generation of potential item bundles that are contextually relevant and tailored to user tastes or requirements.
@@
This comprehensive approach to bundle generation is exemplified by models like (Yang et al., 2024b; Wei et al., 2022), which introduces an encoder-decoder framework capable of generating personalized bundles in a non-autoregressive manner. In contrast, GenRec (Sun et al., 2023b) generates the recommended items one by one to form the next bundle via an autoregressive decoder. DPG (Zheng et al., 2023) integrates a Recurrent Neural Network-based generative model enhanced by a mask layer sensitive to patient conditions, trained via a hybrid approach of maximum likelihood estimation and reinforcement learning to minimize order dependency. DySR (Liu et al., 2021), on the other hand, leverages a GAN-style training strategy to dynamically generate service packages, focusing on the evolution of service representations in a shared semantic space, which is particularly beneficial for addressing cold start challenges and aligning services with user requirements.
@@
In the fashion domain, methods like those proposed by Li et al. (Li et al., 2017) and Han et al. (Han et al., 2017) have traditionally treated fashion bundles as sets or ordered sequences, relying on models such as bidirectional LSTMs for generation. These methods rely heavily on the order of the outfit items. But actually, it is unreasonable to regard an outfit as an ordered sequence, because shuffling items in the outfit should make no difference on its compatibility. Recognizing the limitations of these approaches, Chen et al. (Chen et al., 2019a) introduced POG, an encoder-decoder model that harnesses the power of Transformer architecture to connect user preferences across individual items and outfits, thereby personalizing the fashion outfit generation process. Similar to the approach taken by POG, DiFashion (Xu et al., 2024) leverages the Transformer architecture’s ability to model the relationships between items holistically, and generates outfits by parallel conditional denoising, considering the compatibility and personalization without being constrained by item order.
@@
In trip recommendation, He et al. (He et al., 2019a) presents C-ILP and C-ALNS based on a joint context-aware embedding. Besides, PT-VPMF (Zhao et al., 2021) uses CNNs to capture visual contents in geo-tagged photos, then introduces visual-enhanced probabilistic matrix factorization for personalized recommendation. Due to the outstanding performance of the encoder-decoder model in the TripRec, other methods such as CATHI (Zhou et al., 2019), DeepTrip (Gao et al., 2023b), and CTLTR (Zhou et al., 2022) follow a similar line of reasoning by also employing the RNN or LSTM based model to learn underlying semantic relationships between POIs in the trajectory. CTLTR (Zhou et al., 2022) captures semantic sub-trajectories and corresponding sequential information by using RNN and contrastive learning. Different from these methods, Query2Trip(Wang et al., 2023a) combines sequence generation with Transformer architecture and debiased learning in TripRec modeling to learn more effective representations.

### v2

@@ 3.4. Summary of Existing Discriminative BR Research
(Deng et al., 2020; Chang et al., 2020; Zhao et al., 2022; Wei et al., 2023; Wang et al., 2023a; Ma et al., 2022)
(Liu et al., 2023b; Zhu et al., 2023b; Ke et al., 2023b, a; Ren et al., 2023; Zhang et al., 2022)
@@
(Li et al., 2023d; Lu et al., 2022; Vijaikumar et al., 2021; Zhang et al., 2021b; Liu et al., 2024a; Yu et al., 2022)
- (Chen et al., 2019a; Sun et al., 2022a; Wang et al., 2021b)
Category, Attributes, Genre - (Xu et al., 2023; Yang et al., 2023; Li et al., 2023a, b)
@@ 4.1. Generative BR Task
(22) $\phi^{*}=\arg\max_{\phi}\textstyle\sum_{u\in\mathcal{U}_{\text{train}}}\log p(g_{\text{gen}}(u,b,\mathbf{X}_{M\times N},\mathbf{Y}_{M\times O},\mathbf{Z}_{N\times O},c;\phi)\mid u),$
@@ 4.2. Representation Learning from Item Level
Category Summarization References Category Summarization References
Embedding Models MF-based (Yang et al., 2024b; Zhang et al., 2024a) Feature Engineering Feature Extraction (cost, time, price et al.) (Yu et al., 2015; Fang et al., 2014; Chen et al., 2015; Xie et al., 2010) (Benouaret and Lenne, 2017, 2016; Banerjee et al., 2020; Chen et al., 2016)
DNN-based
(RNN, LSTM et al.)
(Bai et al., 2019; Wei et al., 2022; Han et al., 2017; Chen et al., 2019b; Li et al., 2017; Deng et al., 2023b; Wang et al., 2023b)
(Zheng et al., 2023; Gu et al., 2021; Liu et al., 2023a; Liu and Cao, 2020; Deng et al., 2021, 2023a; Zhou et al., 2019)
(Ma et al., 2024c, b; Xu et al., 2024; Zhang et al., 2024a; Liu et al., 2024b; Gao et al., 2023c)
GNN-based
(GCN, GAT et al.)
@@
(Zheng et al., 2023; Qi et al., 2021; Deng et al., 2021; Li et al., 2023c)
(Ma et al., 2024c; Ding et al., 2023; Yu et al., 2023; Hu et al., 2025)
CF-based
(UserCF, ItemCF)
@@ 4.2.1. Embedding Models
Similar with representation learning of Discriminative BR, both DNN-based and GNN-based models are frequently utilized in Generative BR to learn representations in many studies (Chang et al., 2021; Vijaikumar et al., 2021; Zheng et al., 2023; Yang et al., 2024b; Bai et al., 2019; Chen et al., 2019b).
@@
MF-based. BundleNAT (Yang et al., 2024b) introduces two kinds of signal to learn the needed dependency information, i.e., the preference signal and the item compatibility signal. It employs MF to extract the preference signal from user-item interactions and uses GNNs to capture item compatibility signal from the co-occurrence graph, which can strengthen the intra-relatedness within the bundle. MatTrip (Zhang et al., 2024a) utilizes Funk-SVD for feature completion to extract and complete the user’s interest.
@@
DNN-based. BGN (Bai et al., 2019) uses CNN as an encoder to extract user context information, embedding user based on user purchase history. Conna (Wei et al., 2022) adopts self-attention to learn the representations for mixture types of candidate objects. Users’ interests are usually distributed across multiple aspects and a fine-grained representation learning of user interests from multiple aspects is conducted by (Deng et al., 2023a). However, this approach does not fully consider the comprehensiveness and dynamism of user interests. To address this limitation, a multi-view multi-aspect neural network representation learning method is introduced by (Deng et al., 2023b), which can more accurately model the multi-aspect interests of users and enhance the accuracy of next-basket recommendations. In works such as (Chen et al., 2019b; Han et al., 2017; Li et al., 2017; Ma et al., 2024c, b), items are represented using a multi-modal embedding model based on multimodal inputs, including image, text, category of the items. Recent advancements in large-scale multimodal foundation models, such as BLIP (Li et al., 2022a) and CLAP (Wu et al., 2023b), have paved the way for studies like CLHE (Ma et al., 2024c) and Bundle-LLM (Liu et al., 2024b) to harness these models for extracting the multimodal features of items. These studies have introduced self-attention mechanism to fuse multimodal information and multi-item representation. Furthermore, Bundle-LLM (Liu et al., 2024b) employs CF method to obtain relational features from user-item interactions and bundle-item affiliations, which can enrich the representation of items. CIRP (Ma et al., 2024b) learns high-quality representations of item with multimodal encoders and pre-trained targets including image-text contrastive loss and cross-item contrastive loss. In order to capture the compatibility of items, a masked item prediction task based on a bidirectional Transformer encoder architecture is devised by (Chen et al., 2019b) to capture the item interactions in a bundle. Bi-LSTM is utilized in (Han et al., 2017) to learn item compatibility within a bundle. Similarly, MatTrip (Zhang et al., 2024a) processes user interest and geographic information separately by two independent Bi-LSTM encoders. In the domain of service package generation, the process begins with parsing textual descriptions of services and mashups to distill semantic features. This is achieved through techniques such as Latent Dirichlet Allocation (LDA) and Natural Language Processing (NLP) methods, which convert text into vector representations that capture essential topics and semantics, as discussed in the works (Gu et al., 2021) and (Liu et al., 2023a). Services and mashups are initially represented as static embeddings, using pre-trained word embeddings like Word2Vec or GloVe to encapsulate their inherent attributes and user requirements (Liu and Cao, 2020; Qi et al., 2021).
@@
GNN-based. To address the dynamic nature of services, a dynamic graph neural network is constructed to model the evolving relationships between services. The graph typically includes nodes for services, mashups, categories, and sometime words, with edges defined by various relationships such as co-occurrence, usage frequency, and semantic associations (Liu et al., 2023a). Graph convolution operations, potentially involving GraphTransformer or Graph Convolutional Networks (GCNs), are applied to update the static embeddings by aggregating information from neighboring nodes in the graph. This allows for the dynamic adjustment of service representations in accordance with the network’s topology (Liu and Cao, 2020; Qi et al., 2021). The graph attention mechanism is leveraged in (Vijaikumar et al., 2021) to capture the high-order relationship between users, items and bundles, which can understand the constitution of a bundle better. BGGN (Chang et al., 2021) learns user and item embeddings through GNNs, and uses GCN to learn the complex interactions upon re-constructed bundle graph. Hypergraph convolution operations are performed on hypergraph to learn better and intent-aware representations of hyperedges (Yu et al., 2023; Li et al., 2023c). TOG (Ding et al., 2023) first defines category combinations in outfit as templates, building two bipartite graphs (user-template and user-item), and then employ GNNs to obtain the enhanced representations of users, templates and items. In drug package scenario, DPG (Zheng et al., 2023) models item relationships by constructing item interaction graph based on medical records and domain knowledge. It devises a message-passing neural network to capture drug interaction (synergism, antagonism, etc.) and uses DNNs to obtain the patient embeddings.
@@ 4.2.2. Feature Engineering
In the context of personalized travel package recommendations, Yu et al. leverage CF to dynamically extract user preferences from location-based social network data, constructing user profiles that reflect individual preferences for various POIs (Yu et al., 2015). Similarly, Fang et al. propose a framework that integrates both POIs and travel packages, employing a score inference model to personalize trip recommendations, which takes into account user-based preferences and temporal properties, inferred through CF techniques (Fang et al., 2014). Benouaret et al. further extend the application of CF by designing a package recommendation framework for trip planning. Their system emphasizes the importance of diversity and relevance in package recommendations, using a composite recommendation system inspired by CF to group diverse POIs into packages that align with user preferences and budget constraints (Benouaret and Lenne, 2016). Building on this, Benouaret et al. aim to recommend top-k packages to users, with each package constituted by a set of POIs under specified constraints. They introduce a scoring function that considers user preferences, diversity, and popularity, all of which are integral components of CF-based representation learning (Benouaret and Lenne, 2017).
@@ 4.3. Bundle Generation
Category Techniques References Category Techniques References
@@
(Zheng et al., 2023; Bai et al., 2019; Han et al., 2017; Zhou et al., 2019; Gao et al., 2023c)
(Li et al., 2017; Yu et al., 2023; Hu et al., 2025; Zhou et al., 2022)
Beam Search (Bai et al., 2019; Chang et al., 2021; Yu et al., 2015; Zhang et al., 2024a; Sun et al., 2023b)
@@
(Wei et al., 2022; Yang et al., 2024b; Chen et al., 2019b)
(Sun et al., 2023b; Wang et al., 2023b)
@@ 4.3.1. Search-based
Greedy search and beam search have also been employed to construct personalized travel routes (Gu et al., 2020; Fang et al., 2014; Yu et al., 2015; Zhang et al., 2024a). TRAR (Gu et al., 2020) uses greedy search to generate candidate set by gradually adding POIs to build the optimal travel package that satisfy multiple user-specific constraints. Yu et al. (Yu et al., 2015) utilize a path-tree based beam search method, focusing on iteratively selecting POIs based on the highest transition probabilities to form travel routes. In contrast, the work by Zhang et al. (Zhang et al., 2024a) integrates user preferences, geographic information, and a grid beam search approach within an iterative process to select candidate POIs, generating travel routes that meet individualized needs and constraints. Furthermore, dynamic programming and integer programming have been used to solve search problem of forming optimal travel packages (Chen et al., 2016; Lim et al., 2015).
@@ 4.3.2. Model-based
Search-based methods generate bundles from the candidate set by searching for the optimal combinations. These methods can explore different combinations more intuitively, but face two problems: first, the search process often relies on predefined heuristics, so the generation effect may not be flexible enough in complex scenarios. Second, the search space expands drastically with the increase of the candidate set, leading to a significant increase in computational overhead. At this point, model-based can cope with these problems and better capture complex relationships, generating bundles with higher quality. As introduced in Section 4.2, embedding models and feature extraction techniques are used for distilling features, representing data, and learning compatibility between items. The subsequent phase involves the utilization of these encoded representations by generative models, such as RNNs and GANs. These models are tasked with the generation of potential item bundles that are contextually relevant and aligned with user preferences.
@@
This comprehensive approach to bundle generation is exemplified by models like (Yang et al., 2024b; Wei et al., 2022), which introduce an encoder-decoder framework capable of generating personalized bundles in a non-autoregressive manner. In contrast, GenRec (Sun et al., 2023b) generates the recommended items one by one to form the next bundle via an autoregressive decoder. DPG (Zheng et al., 2023) integrates a RNN-based generative model enhanced by a mask layer sensitive to patient conditions, trained via a hybrid approach of maximum likelihood estimation and reinforcement learning to minimize order dependency. DySR (Liu et al., 2021), on the other hand, leverages a GAN-style training strategy to dynamically generate service packages, focusing on the evolution of service representations in a shared semantic space, which helps address cold start challenges and align services with user requirements.
@@
In the fashion domain, methods like those proposed by Li et al. (Li et al., 2017) and Han et al. (Han et al., 2017) have traditionally treated fashion bundles as sets or ordered sequences, relying on models such as bidirectional LSTMs for generation. These methods rely heavily on the order of the outfit items. But actually, it is unreasonable to regard an outfit as an ordered sequence, because shuffling items in the outfit should make no difference on its compatibility. Recognizing the limitations of these approaches, Chen et al. (Chen et al., 2019b) introduced POG, an encoder-decoder model that harnesses the power of Transformer architecture to connect user preferences across individual items and outfits, thereby personalizing the fashion outfit generation process. Similar to the approach taken by POG, DiFashion (Xu et al., 2024) leverages the Transformer architecture’s ability to model the relationships between items holistically, and generates outfits by parallel conditional denoising, considering the compatibility and personalization without being constrained by item order.
@@
In trip recommendation, He et al. (He et al., 2019a) presents C-ILP and C-ALNS based on a joint context-aware embedding. Besides, PT-VPMF (Zhao et al., 2021) uses CNNs to capture visual contents in geo-tagged photos, then introduces visual-enhanced probabilistic matrix factorization for personalized recommendation. Due to the outstanding performance of the encoder-decoder model in the TripRec, other methods such as CATHI (Zhou et al., 2019), DeepTrip (Gao et al., 2023c), and CTLTR (Zhou et al., 2022) follow a similar line of reasoning by also employing the RNN or LSTM based model to learn underlying semantic relationships between POIs in the trajectory. CTLTR (Zhou et al., 2022) captures semantic sub-trajectories and corresponding sequential information by using RNN and contrastive learning. Different from these methods, Query2Trip(Wang et al., 2023b) combines sequence generation with Transformer architecture and debiased learning in TripRec modeling to learn more effective representations.